Given the following demand function P=3/Q+Q-5

Calculate:

Total Revenue l

Marginal Revenue

Average Revenue



1
Expert's answer
2021-02-26T10:42:30-0500
  1. "TR=P\u00d7Q"

"TR=(\frac{3{}}{Q}+Q⁻⁵)Q"

Total Revenue (TR) = 3+"Q\u207b\u2074"

Marginal Revenue (MR)"=TR'" = "-4Q\u207b\u2075"


Average Revenue (AR) ="\frac{TR{}}{Q}"

AR = "\frac{3+Q^{-4}}{Q}"

AR = "\frac{3}{Q}+Q^{-5}"